plz help me with this plz? iv been up all night trying to figure it out. its 26cm/s=?m/min. plz help
Answer by ptaylor(2198)

plz help me with this plz? iv been up all night trying to figure it out. its
26cm/s=?m/min. plz help

1s=(1/60)min and
1cm=(1/100)m and
26cm=(26/100)m
26cm/s=(26/100)m/(1/60)min multiply numerator and denominator by 60, that is: (60/60)=1
((26)(60)/100)m/((60)(1/60))min=
(1560/100)m/1min=
15.6m/min

26cm/s=15.6m/min

Ck
If it can travel 26cm in 1s, then it can travel (26)(60) or 1560 cm in 60 sec or (1 min)
Now if it can travel 1560cm in 1 min then it can travel (1560/100) or 15.6 meters in 1 min

Hope this helps -----ptaylor